Two skydivers are holding on to each other while f
Answer #1

For first question answer as following

We have v1x = 4.93m/s; v1y = 4.75m/s ; v1z =63.1m/s

            First sky driver mass, m1 = 89.30kg

            Second sky driver mass, m2 =57.7kg

Now

v2x= -v1x x (m1/m2) = - 4.93 x (89.3 /57.7) = -4.93 x1.548 =-7.63 m/s

v2y = -v1y x (m1/m2) = - 4.75 x (89.3 /57.7) = -4.75 x 1.548 =-7.353m/s

v2z = ((m1+m2)63.1-v1z m1)/ m2 = [(147x63.1) – (63.1 x89.30)]/57.7

            =[9275.7 -5634.83] /57.7 =3640.87/57.7 = 63.10m/s

For second question answer as following

               Initial K.E = ½ m v2 = ½ (m1+m2) v2 = ½ (89.3+57.7) (63.1)2

                                    = 0.5 x147x 3981.61 = 292648.34J

Final K.E = ½ (m1v12 +m2v22 )

v12 = v1x2 + v1y2+v1z2 = (4.93)2 +(4.75)2 + (63.1)2 = 24.31+22.56 +3981.61 =4028.48 m2/s2

v22 = v2x2+v2y2+v2z2 = (-7.63)2 + (-7.353)2 + (63.1)2 =58.22 +54.067+3981.61 =4093.897 m2/s2

final K.E = ½ ( 89.3 x 4028.48 + 57.7 x4093.897) = ½ (359743.264 +236217.86)

= 297980.56 J

ΔK.E = 297980.56- 292648.34J =5332.2J